Find specific charge if an cathode ray is passed through cross feilds in terms of (E) (B) and (V).Where E is electric feild, B is magnetic feild and V 8s pot. differene
1
Expert's answer
2020-05-04T12:41:07-0400

As per the given question,

Let the charge of the cathode beam is Q, mass of the charge is m, it is passing in the magnetic field B and electric field E, in the potential difference V. let the radius of the cathode ray is R,

Now, applying the conservation of energy,

"QV=\dfrac{mv^2}{2}------(i)"

Now,

"\Rightarrow QvB=QE"

"\Rightarrow v=\dfrac{E}{B}------(ii)"

now substituting the value of v in equation (i) from equation (ii)

"QV=\dfrac{mE^2}{2B^2}"

"\Rightarrow \dfrac{Q}{m}=\dfrac{E^2}{2B^2V}"